Use completed square form, (×-18)^2 +304, to solve the equation x^2 -36× -20=0 leaving answer in exact (surd) form.
don't know where to start on this so any help would be fab. Thanks

Answer by fractalier(6550)

To solve
x^2 - 36× - 20 = 0
by completing the square goes like this...
We rewrite it this way...
x^2 - 36x + _____ - 20 = 0 + ______
What goes in the blank is found by asking, "What is half of -36 squared?"
Well, half of -36 is -18, and -18 squared is 324.
So now we have
x^2 - 36x + 324 - 20 = 0 + 324
Add twenty to both sides and get
x^2 - 36x + 324 = 344
Now rewrite the left side as a square...
(x - 18)^2 = 344
Now take the square root of both sides...
x - 18 = plus/minus sqrt(344)
and
x = 18 plus or minus sqrt(344) = 18 +/- 2*sqrt(86)